FInd the exact value of tan(A-B) if cos(A)=3/5 where 0< or equal to A
Answer by lwsshak3(11628)

FInd the exact value of tan(A-B) if cos(A)=3/5 where 0< or equal to A ***
cos A=3/5 (working with a (3-4-5) reference right triangle in quadrant I)
sinA=4/5
..
sinB=-5/13 (working with a (5-12-13) reference right triangle in quadrant III)
cosB=-12/13
..
(sin(A-B)=sinAcosB-cosAsinB=4/5*(-12/13)-3/5*(-5/13)=-48/65+15/65=-33/65
cos(A-B)=cosAcosB+sinAsinB=3/5*(-12/13)+4/5*(-5/13)=-36/65-20/65=56/65
tan(A+B)=sin(A+B)/cos(A+B)=-33/56